OpenStudy (falling_in_katt):

Which of the following expressions is a factor of the polynomial x ^ 2 +3/2 x - 1? A. x-1/2 B. x-2 C. x-1 D. x+3/2

OpenStudy (falling_in_katt):

@Luigi0210

OpenStudy (luigi0210):

Is it this? \(\large x^2+\frac{3}{2}x-11\)?

OpenStudy (luigi0210):

Whoops, that's suppose to be 1 not 11

OpenStudy (falling_in_katt):

yes

OpenStudy (luigi0210):

If you know synthetic division it would work great here

OpenStudy (falling_in_katt):

I have no clue how to do that.

OpenStudy (luigi0210):

Well synthetic division is another way of doing long division. All you have to do is take the coefficients of the polynomial and write in an upside down division bar like so:  Since the coefficients of \(x^2+3/2 x -1 \) are 1, 3/2, and -1 On left you put what you are dividing by, so let's try x-1/2 \(x-\frac{1}{2}=0 \) so \(x=\frac{1}{2} \)